The data below are the quiz scores of 10 randomly selected statistics students and the number of hours they studied for the quiz.

Hours, x= 3, 5, 2, 8, 2, 4, 4, 5, 6, 3

Scores, y= 65, 80, 60, 88, 66, 78, 85, 90, 90, 71


What is the best predicted value for the quiz score for a student who studied 2 hours?
Assume that the variables x and y have a significant correlation.

Group of answer choices

67

65

66

64
Answer #1

solution:-

given by the above information

the regression line is

=> y^ = 56.114+5.044 x

explanation for the regression line

X Y XY Χ. Χ 3 65 195 9 5 80 400 25 2 60 120 4 8 88 704 64 2 66 132 4 4 78 312 16 4 85 340 16 5 90 450 25 6 90 540 36 3 71 213

the best predicted value for the quiz score for a student who studied 2 hours

=> y^ = 56.114+5.044 x

=> y^ = 56.114+5.044(2)

=> y^ = 66